OpenStudy (anonymous):

let x = 2tan(theta) to evaluate the integral of: 1/((x^(2)sqrt(x^(2)+4))

OpenStudy (anonymous):

no one did this yet?

OpenStudy (anonymous):

nope lol is it hard or easy for u?

myininaya (myininaya):

\[\int\limits_{}^{}\frac{1}{x^2 \sqrt{x^2+4}} dx\]

myininaya (myininaya):

it tells you what substitution they want you to use

myininaya (myininaya):

\[x=2 \tan(\theta) => dx =2 \sec^2(\theta) d \theta\]

myininaya (myininaya):

\[\int\limits_{}^{}\frac{1}{(2\tan(\theta))^2 \sqrt{(2\tan(\theta))^2+4}} 2 \sec^2( \theta) d \theta\]

OpenStudy (anonymous):

idea is to replace x by \[2\tan(\theta)\] then use the trig identity \[\tan^2(\theta)+1=\sec^2(\theta)\] so clear the radical.

OpenStudy (anonymous):

that is the idea. myininaya is writing it out nicely

OpenStudy (anonymous):

i see, but where do i go from there? i am still confused ]

myininaya (myininaya):

\int\limits_{}^{}\frac{1}{4\tan^2(\theta) \sqrt{4 \tan^2(\theta)+4}} 2 \sec^2(\theta) d \theta \int\limits_{}^{}\frac{1}{\4tan^2(\theta) \sqrt{4}\sqrt{\tan^2(\theta)+1}}2 \sec^2(\theta) d \theta \[\frac{1}{4 \sqrt{4}} \int\limits_{}^{} \frac{1}{\tan^2(\theta) \sqrt{\sec^2(\theta)}} 2 \sec^2(\theta) d \theta\]

myininaya (myininaya):

lol oops

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

\[\int\limits_{}^{}\frac{1}{4\tan^2(\theta) \sqrt{4 \tan^2(\theta)+4}} 2 \sec^2(\theta) d \theta\]

OpenStudy (anonymous):

\[\int\limits_{}^{}\frac{1}{\4tan^2(\theta) \sqrt{4}\sqrt{\tan^2(\theta)+1}}2 \sec^2(\theta) d \theta\]

OpenStudy (anonymous):

you can do the work and i will take the credit.

OpenStudy (anonymous):

hey satellite (i was looking for a soln involving paramatetric eqns) and myininaya can you look at my silo question after your done helping here thx

myininaya (myininaya):

\[\int\limits\limits_{}^{}\frac{1}{4\tan^2(\theta) \sqrt{4 \tan^2(\theta)+4}} 2 \sec^2(\theta) d \theta \] \[\int\limits\limits_{}^{}\frac{1}{4 \tan^2(\theta) \sqrt{4}\sqrt{\tan^2(\theta)+1}}2 \sec^2(\theta) d \theta \] \[\frac{1}{4 \sqrt{4}} \int\limits\limits_{}^{} \frac{1}{\tan^2(\theta) \sqrt{\sec^2(\theta)}} 2 \sec^2(\theta) d \theta \]

myininaya (myininaya):

\[\frac{2}{4 (2)} \int\limits_{}^{}\frac{\sec^2(\theta)}{\tan^2(\theta) \sec(\theta)} d \theta\]

myininaya (myininaya):

\[\frac{ \not{2}}{4 (\not{2})}\int\limits_{}^{}\frac{\sec(\theta)}{\tan^2(\theta)} d \theta\]

OpenStudy (anonymous):

\[\frac{1}{4}\int\csc(\theta)\cot(\theta)d\theta=-\frac{1}{4}\csc(\theta)\]

OpenStudy (anonymous):

and now you have to go back...

myininaya (myininaya):

\[\frac{1}{4}\int\limits_{}^{}\frac{1}{\cos(\theta)} \cdot \frac{\cos^2(\theta)}{\sin^2(\theta)} d \theta\]

OpenStudy (anonymous):

\[\theta =tan^{-1}(\frac{x}{2})\]

myininaya (myininaya):

\[\frac{1}{4} \int\limits_{}^{}\frac{\cos(\theta)}{\sin^2(\theta)} d \theta\]

myininaya (myininaya):

let \[u=\sin(\theta) => du=\cos(\theta) d \theta \] \[\frac{1}{4} \int\limits_{}^{}\frac{1}{u^2} du\]

myininaya (myininaya):

\[\frac{1}{4} \int\limits_{}^{}u^{-2} du= \frac{1}{4} \cdot \frac{u^{-2+1}}{-2+1} +C\]

myininaya (myininaya):

\[\frac{1}{4} \cdot \frac{u^{-1}}{-1}+C=\frac{-1}{4u}+C=\frac{-1}{4\sin(\theta)}+C\]

myininaya (myininaya):

so we need to know what \[\sin(\theta) \] means in terms of x this is where we draw a right triangle because geometry is a blast!

myininaya (myininaya):